# Toward A Causal Framework for Modeling Perception

### v3

@@ 5 Toward a Causal Modeling Framework
How might we describe the process through which a receiver $R\in\mathcal{R}$ constructs its SCM $\mathcal{M}_{R}$? We treat this section as a first approach given the complexity of the task.
@@ 5.1 Key Functionalities
Two processes are central to the framework: categorization and signification. Both have their origins in the social sciences (Loury 2019): categorization entails sorting instances or classes into categories while signification entails representing the social meanings of the categories describing the instances or classes of interest. For our purposes, given the nature of $X\in\mathbf{X}$, categorization implies listing the items (or categories) that describe $X$, while signification implies specifying how the items between $X_{i}$ and $X_{j}$ in $\mathbf{X}$ relate. Notably, in signifying two variables, we would like to specify causal-like statements between them.
For instance, if $X_{1}$ represents SAT scores, then categorizing $X_{1}$ would amount to listing items such as “standardized testing”, “tutoring”, and “performance under pressure”. Similarly, signifying $X_{1}$ relative to zip code $Z$ would amount to stating that “household income” and “school district” are positively related to “standardized testing”. The items “household income” and “school district” are obtained from signifying $Z$.
@@
where $\theta_{i}^{X}$’s are items (or categories) from a that $R$ associates to $X$. It is possible for $\Theta^{R}(X=x)=\emptyset$.
@@
where $p=|\mathbf{X}|$. Each $R\in\mathcal{R}$ comes with its own categorization set. It is implied that $X$ is a variable in the SCM $\mathcal{M}$ used by $R$. For readability, we will write $\Theta^{R}(X)$ and $\vartheta^{R}$.
@@
[…]
Importantly, under structural perception (Definition 4.2) the $\Phi^{R}(X_{i},X_{j})$ is causal statement on the cause-effect pair and the causal effect while under parametrical perception (Definition 4.3) the $\Phi^{R}(X_{i},X_{j})$ is causal statement only on the causal effect. We define the signification set as:
@@
where $p=|\mathbf{X}|$. Each $R\in\mathcal{R}$ comes with its own signification set that is based on its own categorization set. It is implied that $X_{i},X_{j}$ are variables in the SCM $\mathcal{M}$ used by $R$. It is possible that $\Phi^{R}(X_{i},X_{j})=\emptyset$.
Implementation-wise, Definition 5.1 parallels—in a simplified way—knowledge representation formalisms such as ontologies (Gruber 1995) or knowledge graphs (Hogan et al. 2022). We would have $X$ as the class or concept and $\Theta^{R}(X)$ as its attributes and properties. We could, thus, view $\vartheta^{R}$ as semantic model (ontologies) or network of facts (knowledge graph) that $R$ has for $\mathbf{X}$. Similarly, Definition 5.2 relates to methods that argue over or relate on represented knowledge. Works on logic argumentation (Besnard, Cordier, and Moinard 2014) or relational learning (Salimi et al. 2020), among other techniques for argumenting knowledge, could be used here. The idea is to be able to argue how two variables causally relate to each other based on the affinity between the items describing them. Under structural perception, the answer would be structural. Under parametric perception, the answer would be parametric.
We emphasize that both definitions are first approaches to capture how a receiver $R$ derives its SCM $\mathcal{M}_{R}$. Together, $\vartheta^{R}$ and $\varphi^{R}$ motivate the construction of $\mathcal{M}_{R}$. The general idea is that these high-level processes are to Definition 4.1 what the representativeness heuristic is to Definition 3.1.
@@ 5.2 Illustrative Examples
Let us revisit Example 4 under categorization and signification. The goal is to show how receivers $R_{1}$ and $R_{2}$ would construct their SCMs.
@@
(Categorization) In Example 1, scenario (ii), assume the admissions officer $R_{1}$ breaks the tie between the two applicants with the same suitability score $G$ from the ML model $f$ using the applicants’ SAT scores $X_{1}$, high-school GPA $X_{2}$, and address $Z$. Applicants have the same $X_{1}$ and $X_{2}$. We define $\vartheta^{R_{1}}$ of applicants’ $X_{1}$, $X_{2}$, and $Z$ as:
@@
[…]
The first line states the variables $R_{1}$ categorizes; and the second the descriptors for each variable according to $R_{1}$.
For later use, we define a second admissions officer $R_{2}$ that has a similar categorization set but with $\Theta^{R_{2}}(X_{1})=\emptyset$, meaning: $\vartheta^{R_{2}}=\{\Theta^{R_{1}}(X_{2}),\Theta^{R_{1}}(Z)\}$.
@@
(Signification: Structural Perception) Based on $\vartheta^{R_{1}}$ and $\vartheta^{R_{2}}$ from Example 5, we define the corresponding signification sets below. Assume (e.g., due to the college’s bylaws) an admissions officer determines $Y$ using $X_{1}$ and $X_{2}$ only. Hence, $X_{1}\rightarrow Y$ and $X_{2}\rightarrow Y$ are given and $Z\rightarrow Y$ is not allowed; these cause-effect pairs are provided to and shared by $R_{1}$ and $R_{2}$. For $R_{1}$:
@@
[…]
@@
(Signification: Parametrical Perception) Similar to Example 6, consider scenario (ii) from Example 1. Suppose $R_{1}$ and $R_{2}$ now agree on $\mathcal{G}_{R_{1}}$. We formalize the scenario where two admissions officers consider the potential of the same applicant differently. Assume both officers have similar signification sets $\varphi^{R_{1}}$ and $\varphi^{R_{2}}$, except for $Z\rightarrow X_{1}$:
[…]
@@
[…]
@@
In all three examples $Z$ acts as a proxy for socioeconomic background. Our framing here is based on the original example from Kleinberg et al. (2019) in which an admissions officer has to decide on two applicants with the same SAT score but from different neighborhoods. We kept the “controversy” from the original example as we believe it clearly would condition what we view as individual experience.
@@ 5.3 Additional Related Work
Following Section 1.2, we now discuss causal perception within the fair causal ML related work.
@@
Given a SCM $\mathcal{M}$ and DAG $\mathcal{G}$ describing $P(\mathbf{X})$, it is possible to imagine movements between low (or micro) and high (or macro) levels of causal abstraction. Works on causal abstraction (Rubenstein et al. 2017; Beckers, Eberhardt, and Halpern 2019; Massidda et al. 2023) study how causal reasoning materializes and is preserved between these SCM abstractions. These works focus mainly on the notion of consistency, studying whether how we reason on one level is consistent with how we reason on the other level. [^3] See Rubenstein et al. (2017, Definition. 3; Theorem. 6) and Beckers, Eberhardt, and Halpern (2019, Definition 3.1) for details. Intuitively, e.g., intervening individual water particles in a river toward a specific direction should be equivalent to intervening the river flow toward that same direction.
With the processes of categorization (Definition 5.1) and signification (Definition 5.2), we point at the concept of causal abstractions by allowing the receivers to go from the random variable $X$ (high-level representation) to its n descriptors $\{\theta_{1}^{X},\cdots,\theta_{n}^{X}\}$ (low-level representation) and relate these to other variable’s descriptors. We do so by assuming a basic structure for $X$ and without formalizing the causal reasoning that takes place in $\mathcal{M}$ and $\mathcal{G}$ across these two levels of abstraction. Our focus is conceptual and centered on knowledge representation, but we share the goal of formalizing how an agent’s causal reasoning is impacted by the possible abstractions comprised in SCM.
Within fair causal ML, Hu and Kohler-Hausmann (2020) question whether a SCM can capture the meaning behind a sensitive random variables, such as gender, by formalizing it as a single node. The work does not discuss causal abstraction, which should account for this critique, but notably argues for a molecular structure to gender, such that there is gender (the high-level representation) and more granular nodes related to it (the low-level representation) in the DAG that simultaneously define gender as concept. Both categorization and signification are inspired from Hu and Kohler-Hausmann (2020)’s discussion. Recently, Mossé et al. (2025) provide a similar conceptual discussion but based on causal abstractions and focused on discrimination testing. We are unaware of other causal fairness works that discuss different levels of causal reasoning. We add to this line of work by considering how multiple agents might disagree on the abstraction levels when looking at the same (protected) attribute and how that leads to perception.
@@
Given a SCM $\mathcal{M}$ and DAG $\mathcal{G}$, counterfactual fairness (CF) (Kusner et al. 2017) establishes that the observed outcome distribution should be the same as the counterfactual outcome distribution when intervening the protected attribute of an individual. It remains the leading causal ML fairness metric (Makhlouf, Zhioua, and Palamidessi 2020). By definition, these two outcomes belong, respectively, to the observational and counterfactual distributions that are both included in $\mathcal{P}_{\mathbf{X}}$ (4). Extensions to CF, such as path-specific CF (Chiappa 2019), also deal with Equation (4) as these compare the observed distribution with respect to some interventional distribution under $\mathcal{M}$ and $\mathcal{G}$.
CF is always with respect to a single SCM. Hence, in a companion paper to Kusner et al. (2017), Russell et al. (2017) consider the robustness of CF when multiple world views, as in multiple SCMs describing $P(\mathbf{X})$, are used to compute CF. The problem is that different SCMs may lead to different fairness results for the same $P(\mathbf{X})$ (Binkyte et al. 2022). Russell et al. (2017) and later works (e.g., Kilbertus et al. (2019) on hidden confounders) studying the robustness of CF keep the focus on the single SCM: all robustness claims are relative to a SCM. The goal, essentially, is to avoid disagreement by “colliding” these worldviews.
With causal perception, we consider the opposite scenario from Russell et al. (2017) by moving away colliding worlds and expecting disagreement among the receivers. Method-wise, we could indeed measure the robustness of a fairness metric like CF under causal perception (Definition 4.1) given these previous works. However, we diverge from these works conceptually by considering the setting in which disagreement is allowed or even encouraged. To the best of our knowledge, what would amount to robustness in such a setting remains unexplored. Our causal perception definition is our contribution to this new line of work on CF robustness claims. Further, a similar argument applies to cases in which we rely on a SCM $\mathcal{M}$ and DAG $\mathcal{G}$ but may allow multiple inputs or perspectives from stakeholders, such as discrimination testing (Álvarez and Ruggieri 2023, 2025) and synthetic bias data generation (Baumann et al. 2023).
@@ 6 Perception and Fairness
In this section, we explore three fair ML areas that benefit from the proposed formalization of perception.
@@ 6.1 Situated Bias
In previous sections, we avoided using the term “bias” despite the connection between perception as a cognitive phenomenon and biased human decision-making (Kahneman 2011). This choice was motivated by two considerations.. The notion of bias often associated to our chosen view of perception (recall, Section 1.2) does not align directly with the definition commonly used in algorithmic fairness. While the former refers to a departure from rational decision-making as described, e.g., by basic probability laws (Kahneman 2011), the latter refers to “demographic disparities in algorithmic systems that are objectionable for societal reason” (Barocas, Hardt, and Narayanan 2023). The latter interpretation is usually the primary concern in fairness-related discourse. Further, building on this distinction, the notion of bias in algorithmic fairness often carries a negative connotation as one aims to capture objectionable disparities. Now, in principle, there is nothing negative being captured by either Definition 3.1 or Definition 4.1. Both definitions formalize a disagreement in, respectively, probabilistic and causal probabilistic reasoning between two individuals. Thus, we must be more precise by “situating” the bias (Haraway 1988).
The disagreement between receivers $R_{i}$ and $R_{j}$ may lead to bias, but to speak of the kind of bias that drives algorithmic unfairness we must not only have a disagreement but also a preference on which interpretation of the information $\mathbf{X}$ is preferred. Similar to how Haraway (1988) argues that all objective knowledge is based on a partial view of the problem, [^4] Objectivism is a function of what we choose to see (situation), how we choose to see it (location), and from where we choose to see it (position), which are telling of privilege and power relations. our framework could be used to situate bias by defining one receiver’s (causal) probabilistic interpretation of $\mathbf{X}$ as the reference interpretation. The underlying idea is that since individual experience, which we construct through the SCM $\mathcal{M}$, shapes decision-making, then we can use our formalization to explore settings in which, e.g., we cannot reach an agreement on what is fair or want to be explicit on what fairness looks like when reasoning within a context.
@@
In Examples  6 and 7 we could define $R_{1}$ as the representative fair receiver given its reasoning behind $X_{1}$ and $Z$. $R_{1}$ would, thus, represent the desired decision-maker to break potential ties between applicants in terms of the suitability score $G$ of the ML model $f$.

### v4

@@ 5 Toward a Causal Modeling Framework
How can we describe the process through which a receiver $R\in\mathcal{R}$ constructs its SCM $\mathcal{M}_{R}$? We treat this section as a first approach given the complexity of the task.
@@ 5.1 Key Functionalities
Two processes are central to this framework: categorization and signification. While categorization entails sorting instances or classes into categories, signification entails representing the social meanings of the categories describing the instances or classes of interest (Loury 2019). For our purposes, given $X_{i},X_{j}\in\mathbf{X}$ categorization implies listing the items that describe $X_{i}$ and $X_{j}$ while signification implies specifying how these items relate $X_{i}$ and $X_{j}$. Notably, in signifying any two variables in $\mathbf{X}$ we want to specify causal-like statements between them.
To illustrate these processes, for instance, if $X_{1}$ represents SAT scores, then categorizing $X_{1}$ could amount to listing items such as “standardized testing”, “tutoring”, and “performance under pressure”. Similarly, signifying $X_{1}$ relative to zip code $Z$ could amount to stating that “household income” and “school district” are positively related to “standardized testing”. The items “household income” and “school district” are obtained from signifying $Z$.
@@
where $\theta_{i}^{X}$ is the i-th item or category that receiver $R$ associates to $X$. It is possible for $\Theta^{R}(X=x)=\emptyset$.
@@
where $p=|\mathbf{X}|$. Each $R\in\mathcal{R}$ comes with its own categorization set. It is implied that $X$ is a variable in the SCM $\mathcal{M}$ used by $R$. For readability, we write $\Theta^{R}(X)$ and $\vartheta^{R}$.
@@
is one of $X_{i}\overset{\beta}{\rightarrow}X_{j}$ or $X_{j}\overset{\beta}{\rightarrow}X_{i}$, which reads as $X_{i}$ causes $X_{j}$ with effect $\beta$” or vice versa. It is possible for $\Phi^{R}(X_{i},X_{j})=\emptyset$. The idea is that receiver $R$ evaluates the items describing the i-th and j-th variables, derived from its categorization set, and from (the intersection of) these items derives (or not, if the result is $\emptyset$) a causal statement between them that depends on the kind of causal perception.
Under structural causal perception (Definition 4.2) the $\Phi^{R}(X_{i},X_{j})$ is causal statement on the cause-effect pair and the causal effect. Under parametrical causal perception (Definition 4.3) the $\Phi^{R}(X_{i},X_{j})$ is causal statement only on the causal effect. We define the signification set as:
@@
where $p=|\mathbf{X}|$. Each $R\in\mathcal{R}$ comes with its own signification set that is based on its own categorization set. It is implied that $X_{i},X_{j}$ are variables in the SCM $\mathcal{M}$ used by $R$.
Implementation-wise, Definition 5.1 relates to knowledge representation formalisms, such as ontologies (Gruber 1995) or knowledge graphs (Hogan et al. 2022) where $X$ is the class or concept and $\Theta^{R}(X)$ its attributes or properties. Categorization would then amount to building a semantic model (for an ontology) or a network of facts (for a knowledge graph) that $R$ has for $\mathbf{X}$. Similarly, Definition 5.2 relates to methods that argue over represented knowledge, such as logic argumentation (Besnard, Cordier, and Moinard 2014) or relational learning (Salimi et al. 2020), among other techniques for argumenting over knowledge. Signification would then amount to arguing how two variables relate causally to each other based on the affinity between the items describing them. Under structural causal perception the answer would be structural, and under parametrical causal perception the answer would be parametric.
These two definitions are first approaches to capture how the receiver $R$ derives its SCM $\mathcal{M}_{R}$. Together, $\vartheta^{R}$ and $\varphi^{R}$ motivate the construction of $\mathcal{M}_{R}$. The general idea here is that these two high-level processes are to Definition 4.1 what the representativeness heuristic is to Definition 3.1.
@@ 5.2 Illustrative Examples
Let us showcase how two receivers might construct their corresponding SCMs through the processes of categorization and signification, leading to causal perception.
@@
(Categorization) In Example 1, scenario (ii), let us assume that the admissions officer $R_{1}$ breaks the tie between the two applicants with the same suitability score $G$ from the ML model $f$ using the applicants’ SAT scores $X_{1}$, high-school GPA $X_{2}$, and address $Z$. Recall from Example 4 that it is implied that these applicants have the same $X_{1}$ and $X_{2}$. We define $\vartheta^{R_{1}}$ of applicants’ $X_{1}$, $X_{2}$, and $Z$ as:
@@
$=$ $\{\{\text{tutoring, expensive, performative}\},$
$\;\;\{\text{discipline, school funding, potential}\},$
$\;\;\{\text{family income, school district}\}\}.$
The first line states the variables $R_{1}$ categorizes and the second line the descriptors for each variable according to $R_{1}$. Additionally, for later use in the upcoming examples we define a second admissions officer $R_{2}$ that has a similar categorization set to $R_{1}$ but with $\Theta^{R_{2}}(X_{1})=\emptyset$, which means that: $\vartheta^{R_{2}}=\{\Theta^{R_{1}}(X_{2}),\Theta^{R_{1}}(Z)\}$.
@@
(Signification for Structural Perception) Based on $\vartheta^{R_{1}}$ and $\vartheta^{R_{2}}$ from Example 5, we define the corresponding signification sets below. Assume due to the college’s bylaws that an admissions officer determines $Y$ using $X_{1}$ and $X_{2}$ only. Hence, $X_{1}\rightarrow Y$ and $X_{2}\rightarrow Y$ are given and $Z\rightarrow Y$ is not allowed; these cause-effect pairs are provided to and shared by $R_{1}$ and $R_{2}$. For $R_{1}$:
@@
$=$ $\{\{\hat{\Phi}^{R_{1}}(\{\text{family income, school district}\},$
$\;\;\;\{\text{tutoring, expensive, performative}\}),$
$\;\;\hat{\Phi}^{R_{1}}(\{\text{family income, school district}\},$
$\;\;\;\{\text{discipline, school funding, potential}\}),\emptyset,$
$\;\;\hat{\Phi}^{R_{1}}(\{\text{tutoring, expensive, performative}\},$
$\;\;\;\{\text{discipline, school funding, potential}\}\}),$
@@
(Signification for Parametrical Perception) Similar to Example 6, consider scenario (ii) from Example 1. Suppose $R_{1}$ and $R_{2}$ now agree on $\mathcal{G}_{R_{1}}$. We formalize the scenario where two admissions officers consider the potential of the same applicant differently. Assume both officers have similar signification sets $\varphi^{R_{1}}$ and $\varphi^{R_{2}}$, except for $Z\rightarrow X_{1}$:
$\Phi^{R_{1}}(Z,X_{1})$ $=\hat{\Phi}^{R_{1}}(\{\text{family income, school district}\},$
$\;\;\;\;\;\;\;\{\text{tutoring, expensive, performative}\})$
@@
$\Phi^{R_{2}}(Z,X_{1})$ $=\hat{\Phi}^{R_{1}}(\{\text{family income, school district}\},$
$\;\;\;\;\;\;\;\{\text{expensive, performative}\})$
@@
In all three examples the address $Z$ acts as a proxy for socioeconomic background. This is intentional as our analysis is based on the original example from Kleinberg et al. (2019) in which an admissions officer decides between two applicants with the same SAT score but from different neighborhoods. We kept the “controversy” from the original example as it shows how a variable like $Z$ could condition an officer’s decision-making based on how she reasons about $Z$ conditional on her individual experience.
@@ 5.3 Additional Related Work
We extend Section 1.2 by discussing causal perception within the fair causal ML literature.
@@
Given a SCM $\mathcal{M}$ and its DAG $\mathcal{G}$ for $P(\mathbf{X})$, it is possible to imagine movements between low (or micro) and high (or macro) levels of causal abstraction. Works on causal abstraction (Rubenstein et al. 2017; Beckers, Eberhardt, and Halpern 2019; Massidda et al. 2023) study how causal reasoning materializes and is preserved between levels. These works focus mainly on the notion of consistency, studying whether the how we reason on one level is consistent with the how we reason on another level given a SCM $\mathcal{M}$. [^3] See Rubenstein et al. (2017, Definition. 3; Theorem. 6) and Beckers, Eberhardt, and Halpern (2019, Definition 3.1) for details. For example, suppose that we have one SCM describing the flow of water particles in a river and another SCM describing the flow of the river itself: consistency implies that intervening the water particles toward a specific direction should be equivalent to intervening the river toward that same direction.
With the processes of categorization (Definition 5.1) and signification (Definition 5.2), we point at the concept of causal abstraction by allowing the receivers to go from the random variable $X$ (a high-level representation) to its n descriptors $\{\theta_{1}^{X},\cdots,\theta_{n}^{X}\}$ (a low-level representation) and relate these to another variable’s descriptors. We do so by assuming a basic structure for $X$ and without formalizing the causal reasoning that takes place in $\mathcal{M}$ and $\mathcal{G}$ across these two levels of abstraction. Our focus is conceptual and centered on knowledge representation, but we share the goal of formalizing how an agent’s causal reasoning is impacted by the possible abstractions comprised in a SCM.
Hu and Kohler-Hausmann (2020) question whether a SCM can capture the meaning behind a sensitive random variable, such as gender, by formalizing it as a single node in a DAG. Their work does not discuss causal abstraction, which should account for this critique, but notably argues for a molecular structure to sensitive attributes, such that there is gender (the high-level representation) and more granular nodes related to it (the low-level representation) in the DAG that simultaneously define gender as concept. Both categorization and signification are inspired by Hu and Kohler-Hausmann (2020)’s discussion. Mossé et al. (2025) provide a similar conceptual discussion based solely on causal abstractions and focused on discrimination testing. We are unaware of other fair causal ML works that discuss different levels of causal reasoning. We add to this line of work by considering how multiple agents can disagree on the abstraction levels when looking at the same (sensitive) attribute and how that leads to perception.
@@
Given a SCM $\mathcal{M}$ and its DAG $\mathcal{G}$, counterfactual fairness (CF) (Kusner et al. 2017) establishes that the observed outcome should be the same as the counterfactual outcome when intervening the sensitive attribute. It remains the leading causal fairness metric (Makhlouf, Zhioua, and Palamidessi 2020). By definition, these two outcomes belong, respectively, to the observational and counterfactual distributions that are both included in $\mathcal{P}_{\mathbf{X}}$ (4). Extensions to CF, such as path-specific CF (Chiappa 2019), also deal with Equation (4) as these always compare the observed distribution with respect to some interventional distribution given $\mathcal{M}$ and $\mathcal{G}$.
CF is with respect to a single SCM. In a companion paper to Kusner et al. (2017), Russell et al. (2017) consider the robustness of CF when multiple world views, as in multiple SCMs describing $P(\mathbf{X})$, are used to compute CF. Although multiple SCMs can be problematic as different SCMs lead to different CF results for the same $P(\mathbf{X})$ (Binkyte et al. 2022), Russell et al. (2017) and later works (see, e.g., Kilbertus et al. (2019) on hidden confounders) study the robustness of CF with a focus on the single SCM: all robustness claims are relative to a SCM of interest. The goal is to avoid disagreement by “colliding” the worldviews into one.
With causal perception we consider the opposite scenario from Russell et al. (2017). We move away from colliding worlds by expecting disagreement among the receivers. Method-wise, we could indeed measure the robustness of a fairness metric like CF under causal perception (Definition 4.1) given these previous works. However, we diverge from them conceptually by considering the setting in which disagreement is allowed or even encouraged. To the best of our knowledge, what would amount to CF robustness in our setting remains unexplored. Our causal perception definition is our contribution to this new line of work on alternative CF robustness claims. A similar argument applies to cases in which we rely on a SCM $\mathcal{M}$ and DAG $\mathcal{G}$ but may allow multiple perspectives from stakeholders, such as discrimination testing (Álvarez and Ruggieri 2023, 2025) and synthetic biased data generation (Baumann et al. 2023).
@@ 6 Perception and Fairness
Finally, we explore three fair ML areas that benefit from the proposed formalization of (causal) perception.
@@ 6.1 Situated Bias
In previous sections, we avoided the term “bias” despite the connection between perception as a cognitive phenomenon and biased human decision-making (Kahneman 2011). This choice was motivated by two considerations. First, the notion of bias often associated to our chosen view of perception (recall, Section 1.2) does not align directly with the definition commonly used in algorithmic fairness. While the former refers to a departure from rational decision-making as described by basic probability laws (Kahneman 2011), the latter refers to “demographic disparities in algorithmic systems that are objectionable for societal reason” (Barocas, Hardt, and Narayanan 2023). The latter interpretation is usually the primary concern in fairness-related discourse. Second, building on this distinction, the notion of bias in algorithmic fairness often carries a negative connotation as one aims to capture objectionable disparities. There is nothing, in principle, negative being captured by Definitions 3.1 and 4.1. Both definitions simply formalize a disagreement in probabilistic and causal probabilistic reasoning between two individuals. We must, thus, be more precise by “situating” the bias (Haraway 1988).
The disagreement between receivers $R_{i}$ and $R_{j}$ may lead to bias, but to speak of the kind of bias that drives algorithmic unfairness we must not only have a disagreement but also a preference on which interpretation of the information $\mathbf{X}$ is preferred. Similar to how Haraway (1988) argues that all objective knowledge is based on a partial view of the problem, [^4] Objectivism is a function of what we choose to see (situation), how we choose to see it (location), and from where we choose to see it (position), which are telling of privilege and power relations. our framework can be used to situate bias by defining one receiver’s (causal) probabilistic interpretation of $\mathbf{X}$ as the reference interpretation. The underlying idea is that since individual experience, which we construct through a SCM $\mathcal{M}$, shapes decision-making, then we can use our formalization to explore settings in which we cannot reach an agreement on what is fair or want to be explicit on what fairness looks like within a specific context.
@@
In Examples  6 and 7 we ca define $R_{1}$ as the representative fair receiver given its reasoning behind $X_{1}$ and $Z$. $R_{1}$ would then represent the desired decision-maker for breaking potential ties between applicants in terms of the suitability score $G$ of the ML model $f$.